OpenStudy (anonymous):

sin(2arcsin(-1/3)) please explain how to get this without the use of a calculator

OpenStudy (amistre64):

table of sines id suppose

OpenStudy (amistre64):

or draw a picture

OpenStudy (amistre64):

since sin is defined as over/hyp; we have a triangle that is 1 over and 3 hyp but somewhere on the underside of a circle

OpenStudy (amistre64):

doubling that "angle" will give us the values of x and y for the outter sin

OpenStudy (amistre64):



OpenStudy (amistre64):

since the angle is going to have the same absolute value; it might be simpler to draw it in a normal position tho

OpenStudy (amistre64):

sin(2t) = 2sin(t)cos(t) right?

OpenStudy (amistre64):

sin(t) = -1/3 cos(t) = sqrt(1-sin^2(t)) = sqrt(1 - 1/9) = sqrt(8/9) = 2sqrt(2)/3 2*-1/3*2sqrt(2)/3 = -4sqrt(2)/9 i believe

OpenStudy (amistre64):